OpenStudy (anonymous):

Modulus & Conjugate of :

OpenStudy (anonymous):

\[\frac{ \sqrt{3}+i }{\sqrt{3}-i }\]

OpenStudy (anonymous):

what is the modulus and conjugate of above ^

OpenStudy (anonymous):

√3+i/ √3+i

OpenStudy (anonymous):

thats the conjugate

OpenStudy (anonymous):

and modulus?

OpenStudy (anonymous):

sorry.. never heard of it.. i only know conjugate

OpenStudy (callisto):

\[\frac{ \sqrt{3}+i }{\sqrt{3}-i }\]\[=\frac{ \sqrt{3}+i }{\sqrt{3}-i }\times \frac{ \sqrt{3}+i }{\sqrt{3}+i }\]\[=\frac{ (\sqrt{3}+i)^2 }{3+1 }\]\[=\frac{ (3-1)+2\sqrt{3}i }{4 }\]\[=\frac{1+ \sqrt{3}i }{2 }\] Find the modulus and conjugate from here.

OpenStudy (anonymous):

how am i going to find a modulus of fraction

OpenStudy (callisto):

How do you find modulus of a complex number z = a + bi?